FT-IR SPECTROSCOPY OF MATRIX ISOLATED PHENYLPEROXYL RADICALS

\underline{ALAN T. MACCARONE}, G.B. ELLISON, Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309-0215; MARK R. NIMLOS, Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309-0215 and National Renewable Energy Laboratory, 1617 Cole Boulevard, Golden, Colorado 80401.

2A′′ phenylperoxyl radicals, C_{6}H_{5}O_{2}, have been generated in a cryogenic Ar matrix. In a two pulse experiment, phenyl radicals, C_{6}H_{5}, are generated via pyrolysis of phenyliodide in Ar through a hyperthermal SiC nozzle and deposited on a CsI window at 20 K; a second valve deposits a layer of O_{2} in Ar on top of the phenyl radicals. The process is repeated and phenylperoxyl radicals are formed in the matrix. The matrix is analyzed with an IR spectrometer to identify vibrational frequencies of the radical.